find the slope and the y intercept of the line
9x = 2y + 8 ? what is the slope

2)FIND THE SLOPE- (3,4) AND (9, -10)? SLOPE M IS WHAT
3)DETERMINE WHETHER THE GRAPH OF THE TWO EQUATION ARE PERPENDICULAR
Y=2X- 3
6Y= 5-X
3)FIND AN EQUATION OF THE LINE CONTAINING THE GIVEN PAIR OF POINTS
(-2, -8) AND (-9, -3)
WHAT IEQUATION OF THE LINE IS Y

Answer by stanbon(75887)

find the slope and the y intercept of the line
9x = 2y + 8 ? what is the slope
Solve for Y:
y = (9/2)x-4
slope = (9/2)
----------------
2)FIND THE SLOPE- (3,4) AND (9, -10)? SLOPE M IS WHAT
slope = (-10-4)/(9-3) = -14/4 = -7/2
---------------------------
3)DETERMINE WHETHER THE GRAPH OF THE TWO EQUATION ARE PERPENDICULAR
Y=2X- 3
slope = 2
------------
6Y= 5-X
slope = (-1/6)
----------------------
The product of the two slopes is not -1, so the lines are not perpendicular.
---------------------------------------
3)FIND AN EQUATION OF THE LINE CONTAINING THE GIVEN PAIR OF POINTS
(-2, -8) AND (-9, -3)
WHAT IEQUATION OF THE LINE IS Y
slope = (-3--8)/(-9--2) = 5/-7 = -5/7
Y= mx + b and you know y,x and m: solve for "b" as follows:
-3 = (-5/7)(-9)+b
-3 = (45/7) + b
-21/7 = 45/7 + b
b = -66/7
---------------
EQUATION:
y = (-5/7)x -(66/7)
==========================
Cheers,
Stan H.